OpenStudy (inowalst):

Let f(x) = 7x - 13. Find f^-1(x).

OpenStudy (amorfide):

to get the inverse you want to make x the subject let's let f(x)=y so we get y=7x-13 y+13=7x divide by 7 (y+13)/7=x then you replace y with x, and let x be f^-1(x) to get (x+13)/7=f^-1(x)

OpenStudy (inowalst):

Oh okay. Thank you.

OpenStudy (theraggedydoctor):

dot.